# On the cardinality of measures of maximal relative entropy for smooth skew products

### v1

@@ 1. Introduction and Main results
In the setting of smooth diffeomorphisms, Newhouse proved that $\mathcal{C}^{\infty}$ diffeomorphisms on compact smooth manifolds always admit a measure of maximal entropy [34]. Buzzi later conjectured that for surface diffeomorphisms with positive topological entropy, there are at most countably many ergodic measures of maximal entropy [15, Conjecture 1]. Sarig’s groundbreaking symbolic coding for $C^{1+\alpha}$ surface diffeomorphisms confirmed this countability and also settled a conjecture of Katok on periodic-orbit growth [40]. This result was recently sharpened by Buzzi–Crovisier–Sarig, proving that $\mathcal{C}^{\infty}$ surface diffeomorphisms with positive entropy have only finitely many ergodic measures of maximal entropy, and exactly one in the transitive case, [17], which answered another of Buzzi’s conjectures [15, Conjecture 2].
In higher dimensions, Ben Ovadia extended Sarig’s coding to non-uniformly hyperbolic diffeomorphisms on arbitrary manifolds, yielding countability results for hyperbolic measures of maximal entropy [9]. More recently, Buzzi–Crovisier–Sarig established the strong positive recurrence (SPR) property, and via SPR obtained strong statistical properties for the associated invariant measures, including exponential decay of correlations and an almost-sure invariance principle [18]. Related results hold for certain two-dimensional fibre skew products with hyperbolic base dynamics, due to Marín–Poletti–Veiga, and for flows [31, 16, 30].
@@ 1.1. Measures of maximal relative entropy
This paper concerns the amount of $\mathbb{P}$-relative entropy that $\Theta$ may admit. Below, we define the metric entropy via the Brin–Katok formula [14] since it is the shortest way to define it (see also [46, Chapter 4]). We also define the concept of $\mathbb{P}$-relative entropy for skew products via an extension of the Brin–Katok formula (see [50]).
@@
We recall that if $h_{\mathbb{P}}(\theta)<\infty$, which will always be the case in this paper, we have that $h_{\mu}(\Theta\mid\mathbb{P})=h_{\mu}(\Theta)-h_{\mathbb{P}}(\theta)$ (see [24, Chapter 5, below equation 1.1.5]).
@@ 1.2. Main results
Hence, measures of maximal $\mathbb{P}$-relative entropy for $\Theta$ correspond exactly to measures of maximal entropy for $T$. In particular, Theorem A yields the usual countability statement for ergodic hyperbolic measures of maximal entropy in the non–skew-product setting in conformance with [9, Theorem 0.5].
@@
In dimension two, it may be possible to strengthen the conclusion of “at most countably many measures of maximal $\mathbb{P}$-relative entropy” of Corollary B to “at most finitely many measures of maximal $\mathbb{P}$-relative entropy”, in analogy with the classical non-skew-product case [17, Corollary 1.4]. Another natural future direction is to derive a criterion for uniqueness. In [17] (see also [18]) it is shown that each homoclinic class supports a unique measure of maximal entropy. In particular, any smooth, transitive surface system with positive entropy has a unique measure of maximal entropy. The analogous statement for measures of relative maximal entropy remains unclear.
@@ 1.3. An application of Corollary B
Sinai conjectured that for $k$ sufficiently large, $T$ has positive metric entropy with respect to Lebesgue measure [42, Page 144]. This conjecture remains open. In contrast, much more is known about measures of maximal entropy: Obata showed that, for large $k$, the standard map admits a unique measure of maximal entropy; moreover, such a measure has positive metric entropy [35]. However, less is known in the random setting. Our results show that there exists at least one and, at most, countably many ergodic relative measures of maximal entropy for certain classes of random perturbations of the standard map.
@@ 2. Hyperbolic measures and Borel homoclinic classes
In this section, we collect standard facts from hyperbolic dynamics and introduce the concept of Borel homoclinic classes (see [18, Definition 2.7]). Fix a compact smooth manifold $E$ and a $\mathcal{C}^{1+\alpha}$ diffeomorphism $\Theta:E\to E$, for some $\alpha>0$. Unless stated otherwise, all statements in this section are made under this standing assumption. For the purposes of this paper, the set $E$ can be thought of as $\Omega\times M$ as in Theorem A. Our exposition closely follows [18, Section 2].
@@ 2.1. Hyperbolic measures
Finally, we recall that if (i) holds but (ii) fails, then $\mu$ is supported on a periodic orbit and hence has zero metric entropy [26, Corollary S.5.2] (see also the proof of [25, Theorem 4.2]).
@@
Below we state Pesin’s stable manifold theorem, whose proof can be found at [38] (see also [18, Theorem 2.2]).
@@
We recall that the choice of $\varepsilon>0$ in Theorem 2.3 does not affect the definition of the global stable and unstable manifolds [8, §8.2] (see also [18, Section 2]).
@@ 3. Topological Markov shifts
equipped with the metric $d(v,u)=\min\{\exp(-\min\{|n|;v_{n}\neq u_{n}\})\}$, and $\sigma:\Sigma\to\Sigma$ is the left shift map $\sigma((v_{i})_{i\in\mathbb{Z}})=(v_{i+1})_{i+1}$. Given a topological Markov shift $\Sigma$ associated to a graph $\mathrm{G}$ we denote as $V(\Sigma)$ the alphabet of $\Sigma$ and we define the set
$\Sigma^{\#}=\left\{(v_{i})_{i\in\mathbb{Z}}\in\Sigma;\,\exists u^{\prime},v^{\prime}\in V(G)\ \text{s.t. }\,\#\{i\in\mathbb{Z}_{+};v_{i}=v^{\prime}\}=\#\{i\in\mathbb{Z}_{-};v_{i}=u^{\prime}\}=\infty\right\}.$ (2)
@@
Let $\theta:\Omega\to\Omega$ be a homeomorphism. We say that $\theta$ admits a Bowen partition if there exist a family of sets $\mathscr{P}\subset\mathcal{B}(\Omega)$ such that
@@
    is a well-defined, Hölder continuous and each $\pi^{-1}(\omega)\leq k_{\mathscr{\mathscr{P}}}$ for some $k_{\mathscr{P}}\in\mathbb{P}$; and
@@
    the function $\pi$ is surjective over the non-wandering set $NW(\theta)$ and there exists $k<\infty$ such that $\pi^{-1}(\omega)\leq k$ for every $\omega\in NW(\theta)$.
Observe that under the above assumptions imply that $\pi\circ\sigma=\theta\circ\pi.$
Theorem 3.3 ([13], [37, Appendix III]).
@@
The next statement, taken from [18, Theorem 10.5 and Section 10], is a higher-dimensional generalisation of Sarig’s celebrated theorem [40, Theorem 1.3] (see also [9, Theorem 0.1]). We mention that the original result presented in [18] is more general; however, the presented version is enough for our purposes.
Theorem 3.4 ([18, Theorem 10.5 and Section 10]).
@@
    $\pi:(R_{i})_{i\in\mathbb{Z}}\in\Sigma\mapsto\bigcap_{i\in\mathbb{Z}}\Theta^{-1}(R_{i})\in\mathcal{X}$
    is well defined and Hölder continuous. Moreover, $f\circ\pi=\pi\circ\sigma$ and $\pi:\Sigma^{\#}\to\mathcal{X}$ is finite-to-one (see Definition 3.1).
@@
Diagram 1: Commutative diagram of the maps $\Theta,\pi,\pi_{s},\pi_{b},\mathrm{proj}_{\Sigma},\mathrm{proj}_{\Sigma_{b}}$, and $\mathrm{proj}_{\Omega}$. We note that $\Sigma_{s}\subsetneq\Sigma_{b}\times\Sigma,$ in fact, $\Sigma_{s}=\{\underline{(P,R)};\underline{P}\in\Sigma_{s},\,\underline{R}\in\Sigma\ \text{and }\pi_{b}(\underline{P})=\mathrm{proj}_{\Omega}\circ\pi(\underline{R})\}$.
@@
$\mathscr{V}_{s}:=\{(P,R);\,R\in\mathscr{R},P\in\mathscr{P}\ \text{and }R\cap(P\times M)\neq\emptyset\},$
@@
$\mathscr{E}_{s}:=\left\{(P,R)\to(P^{\prime},R^{\prime});\,\Theta^{-1}(R^{\prime})\cap R\neq\emptyset\ \text{and }P^{\prime}\cap\theta^{-1}(P)\neq\emptyset\right\}.$
@@
$\pi_{s}:((P_{i},R_{i}))_{i\in\mathbb{Z}}\in\Sigma_{s}\mapsto\bigcap_{i\in\mathbb{Z}}R_{i}\in\Omega\times M.$
@@
$\mu(\pi_{s}(\Sigma))\geq\mu(\pi_{s}(\Sigma^{\#}))\geq\mu(\pi(\Sigma^{\#}))=1$
@@
First of all, observe that if $(\omega,x)=\pi(\underline{R})$. For each $i\in\mathbb{Z}$ let $P_{i}\in\mathscr{P}$ such that $\theta^{i}(\omega)\in P_{i}$. It is clear that $\underline{P}=(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}$. Moreover, $\Theta^{i}(\omega,x)=(\theta\omega,T_{\omega}(x))\in R_{i}$ for each $i\in\mathbb{Z}$. In this way,
$R_{i}\cap(P_{i}\times M)\neq\emptyset,\,\Theta^{-1}(R_{i+1})\cap R_{i}\neq\emptyset\ \text{and}\ \theta^{-1}(P_{i+1})\cap P_{i}\neq\emptyset\ \text{for every }i\in\mathbb{Z}.$
@@
$=\#\left\{\underline{(P,R)}\in\Sigma_{s};\theta^{i}\omega\in P_{i}\ \text{and }R_{i}\cap(P_{i}\times M)\neq\emptyset\ \text{for each }i\in\mathbb{Z}\right\}$
@@
Hence $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ has uniformly bounded fibres and is therefore finite-to-one. Independently, since $\pi_{b}:\Sigma_{b}\to\Omega$ is obtained from a Bowen partition for $\theta$, it is finite-to-one on $NW(\theta)$. This proves item (3).
@@
$\Theta^{i}(\omega,x)=(\theta^{i}(\omega),T_{\omega}^{i}(x))\in R_{i}\ \text{and }R_{i}\cap(P_{i}\times M)\neq\emptyset\text{ for every }i\in\mathbb{N}.$
In particular $\theta^{i}(\omega)\in P_{i}$ for each $i\in\mathbb{Z}$ and therefore $\pi_{b}(\underline{P})=\omega.$ Hence
@@
$\mathscr{V}^{\prime}=\{P\in\mathcal{P}^{\prime};\text{there exists }R\in\mathscr{R}\ \text{such that }(P\times M)\cap R\neq\emptyset\},$
@@
$\mathscr{E}^{\prime}=\left\{P\to P^{\prime}\;\middle|\;\begin{aligned} &P,P^{\prime}\in\mathscr{V}^{\prime},\ \theta^{-1}(P^{\prime})\cap P\neq\emptyset\ \text{and there exist }R,R^{\prime}\in\mathscr{R},\\ &\ \text{such that }\Theta^{-1}(R^{\prime}\cap(P^{\prime}\times M))\cap(R\cap(P\times M))\neq\emptyset\end{aligned}\right\}.$
@@
Since $\mathcal{X}$ is a Borel homoclinic class, recall that $\Sigma$ is an irreducible topological Markov shift. Therefore, there exists $\underline{y}\in\Sigma$ such that $\Sigma=\overline{\{\sigma^{i}\left(\underline{y}\right)\}_{i\in\mathbb{Z}}}.$ Take $\underline{x}\in\Sigma_{s}$ such that $\mathrm{proj}_{\Sigma_{s}}(\underline{x})=\underline{y}$. Let $\mathscr{G}^{\prime}_{s}\subset\mathscr{G}_{s}$ the maximal irreducible component of $\mathscr{G}_{s}$ containing $x_{0}$ (the 0-th element of $\underline{x}$) and define

### v2

@@ 1. Introduction and Main results
In the setting of smooth diffeomorphisms, Newhouse proved that $\mathcal{C}^{\infty}$ diffeomorphisms on compact smooth manifolds always admit a measure of maximal entropy [39]. Buzzi later conjectured that for surface diffeomorphisms with positive topological entropy, there are at most countably many ergodic measures of maximal entropy [16, Conjecture 1]. Sarig’s groundbreaking symbolic coding for $C^{1+\alpha}$ surface diffeomorphisms confirmed this countability and also settled a conjecture of Katok on periodic-orbit growth [45]. This result was recently sharpened by Buzzi–Crovisier–Sarig, proving that $\mathcal{C}^{\infty}$ surface diffeomorphisms with positive entropy have only finitely many ergodic measures of maximal entropy, and exactly one in the transitive case, [18], which answered another of Buzzi’s conjectures [16, Conjecture 2].
In higher dimensions, Ben Ovadia extended Sarig’s coding to non-uniformly hyperbolic diffeomorphisms on arbitrary manifolds, yielding countability results for hyperbolic measures of maximal entropy [10]. More recently, Buzzi–Crovisier–Sarig established the strong positive recurrence (SPR) property, and via SPR obtained strong statistical properties for the associated invariant measures, including exponential decay of correlations and an almost-sure invariance principle [19]. Related results are known for two-dimensional fibre skew products over hyperbolic bases; see Marín–Poletti–Veiga [37]. Analogous results are available for flows [35, 17, 33] and for discontinuous or maps with singularities [32, 31, 4, 34].
@@ 1.1. Measures of maximal relative entropy
This paper concerns the amount of $\mathbb{P}$-relative entropy that $\Theta$ may admit. Below, we define the metric entropy via the Brin–Katok formula [15] since it is the shortest way to define it (see also [51, Chapter 4]). We also define the concept of $\mathbb{P}$-relative entropy for skew products via an extension of the Brin–Katok formula (see [55]).
@@
We recall that if $h_{\mathbb{P}}(\theta)<\infty$, which will always be the case in this paper, we have that $h_{\mu}(\Theta\mid\mathbb{P})=h_{\mu}(\Theta)-h_{\mathbb{P}}(\theta)$ (see [25, Chapter 5, below equation 1.1.5]).
@@ 1.2. Main results
Hence, measures of maximal $\mathbb{P}$-relative entropy for $\Theta$ correspond exactly to measures of maximal entropy for $T$. In particular, Theorem A yields the usual countability statement for ergodic hyperbolic measures of maximal entropy in the non–skew-product setting in conformance with [10, Theorem 0.5].
@@
In dimension two, it may be possible to strengthen the conclusion of “at most countably many measures of maximal $\mathbb{P}$-relative entropy” of Corollary B to “at most finitely many measures of maximal $\mathbb{P}$-relative entropy”, in analogy with the classical non-skew-product case [18, Corollary 1.4]. Another natural future direction is to derive a criterion for uniqueness. In [18] (see also [19]) it is shown that each homoclinic class supports a unique measure of maximal entropy. In particular, any smooth, transitive surface system with positive entropy has a unique measure of maximal entropy. The analogous statement for measures of relative maximal entropy remains unclear.
@@ 1.3. An application of Corollary B
Sinai conjectured that for $k$ sufficiently large, $T$ has positive metric entropy with respect to Lebesgue measure [47, Page 144]. This conjecture remains open. In contrast, much more is known about measures of maximal entropy: Obata showed that, for large $k$, the standard map admits a unique measure of maximal entropy; moreover, such a measure has positive metric entropy [40]. However, less is known in the random setting. Our results show that there exists at least one and, at most, countably many ergodic relative measures of maximal entropy for certain classes of random perturbations of the standard map.
@@ 2. Hyperbolic measures and Borel homoclinic classes
In this section, we collect standard facts from hyperbolic dynamics and introduce the concept of Borel homoclinic classes (see [19, Definition 2.7]). Fix a compact smooth manifold $E$ and a $\mathcal{C}^{1+\alpha}$ diffeomorphism $\Theta:E\to E$, for some $\alpha>0$. Unless stated otherwise, all statements in this section are made under this standing assumption. For the purposes of this paper, the set $E$ can be thought of as $\Omega\times M$ as in Theorem A. Our exposition closely follows [19, Section 2].
@@ 2.1. Hyperbolic measures
Finally, we recall that if (i) holds but (ii) fails, then $\mu$ is supported on a periodic orbit and hence has zero metric entropy [27, Corollary S.5.2] (see also the proof of [26, Theorem 4.2]).
@@
Below we state Pesin’s stable manifold theorem, whose proof can be found at [43] (see also [19, Theorem 2.2]).
@@
We recall that the choice of $\varepsilon>0$ in Theorem 2.3 does not affect the definition of the global stable and unstable manifolds [9, §8.2] (see also [19, Section 2]).
@@ 3. Topological Markov shifts
equipped with the metric $d(v,u)=\min\{\exp(-\min\{|n|;v_{n}\neq u_{n}\})\}$, and $\sigma:\Sigma\to\Sigma$ is the left shift map $\sigma((v_{i})_{i\in\mathbb{Z}})=(v_{i+1})_{i+1}$. Given a topological Markov shift $\Sigma$ associated to a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ we define the set
$\Sigma^{\#}:=\left\{(v_{i})_{i\in\mathbb{Z}}\in\Sigma;\,\exists u^{\prime},v^{\prime}\in\mathscr{V}\ \text{s.t. }\,\#\{i\in\mathbb{Z}_{+};v_{i}=v^{\prime}\}=\#\{i\in\mathbb{Z}_{-};v_{i}=u^{\prime}\}=\infty\right\}.$ (2)
@@
Let $\theta:\Omega\to\Omega$ be a homeomorphism. We say that $\theta$ admits a Bowen partition if there exist a family of compact sets $\mathscr{P}\subset\mathcal{B}(\Omega)$ such that
@@
    is well-defined and Hölder continuous; and
@@
    the function $\pi$ is surjective over the non-wandering set $NW(\theta)$ and $\pi^{-1}(\omega)\leq k_{\mathscr{P}}$ for some $k_{\mathscr{P}}\in\mathbb{N}$ for every $\omega\in NW(\theta)$.
Observe that under the above assumptions imply that $\pi_{b}\circ\sigma=\theta\circ\pi_{b}.$
Theorem 3.3 ([14], [42, Appendix III]).
@@
The next statement, taken from [19, Theorem 10.5 and Section 10], is a higher-dimensional generalisation of Sarig’s celebrated theorem [45, Theorem 1.3] (see also [10, Theorem 0.1]). We mention that the original result presented in [19] is more general; however, the presented version is enough for our purposes.
Theorem 3.4 ([19, Theorem 10.5 and Section 10]).
@@
    $\pi:(R_{i})_{i\in\mathbb{N}}\in\Sigma\mapsto\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\in\mathcal{X}$
    is well defined and Hölder continuous. Moreover, $\Theta\circ\pi=\pi\circ\sigma$ and $\pi:\Sigma^{\#}\to\mathcal{X}$ is finite-to-one (see Definition 3.1).
@@
Diagram 1: Commutative diagram of the maps $\Theta,\pi,\pi_{s},\pi_{b},\mathrm{proj}_{\Sigma},\mathrm{proj}_{\Sigma_{b}}$, and $\mathrm{proj}_{\Omega}$. We note that $\Sigma_{s}\subsetneq\Sigma_{b}\times\Sigma,$ in fact, $\Sigma_{s}\subset\{\underline{(P,R)};\underline{P}\in\Sigma_{b},\,\underline{R}\in\Sigma\ \text{and }\pi_{b}(\underline{P})=\mathrm{proj}_{\Omega}\circ\pi(\underline{R})\}$.
@@
$\mathscr{V}_{s}:=\{(P,R);\,R\in\mathscr{R},P\in\mathscr{P}\ \text{and }\overline{R}\cap(P\times M)\neq\emptyset\},$
@@
$\mathscr{E}_{s}=\left\{(P,R)\to(P^{\prime},R^{\prime})\;\middle|\;\begin{aligned} &(P,R),(P^{\prime},R^{\prime})\in\mathscr{V}_{s},\ P^{\prime}\cap\theta^{-1}(P)\neq\emptyset\\ &\text{and }\Theta^{-1}(R^{\prime})\cap R\neq\emptyset\end{aligned}\right\}.$
@@
$\pi_{s}:((P_{i},R_{i}))_{i\in\mathbb{Z}}\in\Sigma_{s}\mapsto\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\in\Omega\times M.$
Observe that $\pi_{s}((P_{i},R_{i})_{i\in\mathbb{Z}})=\pi((R_{i})_{i\in\mathbb{Z}}).$
@@
$\mu(\pi_{s}(\Sigma_{s}))\geq\mu(\pi_{s}(\Sigma_{s}^{\#}))\geq\mu(\pi(\Sigma^{\#}))=1$
@@
Let $(\omega,x)\in\Omega\times M$ such that $\pi(\underline{R})=(\omega,x)$. From the definition of $\pi$ we obtain that
$(\omega,x)\in\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\ \text{for each }n\in\mathbb{N}.$
In particular, $\Theta^{i}(\omega,x)=(\theta^{i}\omega,T_{\omega}^{i}(x))\in\overline{R_{i}}$ for each $i\in\mathbb{N}.$ For each $i\in\mathbb{N}$, let $P_{i}\in\mathscr{P}$ such that $\theta^{i}\omega\in P_{i}.$ Therefore, we obtain that
$\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset\ \text{for each }i\in\mathbb{N}.$
In this way,
$\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset,\,\Theta^{-1}(R_{i+1})\cap R_{i}\neq\emptyset\ \text{and}\ \theta^{-1}(P_{i+1})\cap P_{i}\neq\emptyset\ \text{for every }i\in\mathbb{Z}.$
@@
$=\#\left\{\underline{(P,R)}\in\Sigma_{s};\theta^{i}\omega\in P_{i}\ \text{and }\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset\ \text{for each }i\in\mathbb{Z}\right\}$
@@
Hence $\pi_{s}:\Sigma_{s}^{\#}\to\mathcal{X}$ is finite-to-one. Independently, since $\pi_{b}:\Sigma_{b}\to\Omega$ is obtained from a Bowen partition for $\theta$, it is finite-to-one on $NW(\theta)$. This proves item (3).
@@
$\overline{R_{i}}\cap(P_{i}\times M)\neq\emptyset\ \text{and }\theta^{i}(\omega)\in P_{i}\text{ for every }i\in\mathbb{N}.$
In particular, we obtain that $\pi_{b}(\underline{P})=\omega.$
Hence
@@
$\mathscr{V}^{\prime}=\{P\in\mathscr{P};\text{there exists }R\in\mathscr{R}\ \text{such that }\ \overline{R}\cap(P\times M)\neq\emptyset\},$
@@
$\mathscr{E}^{\prime}=\left\{P\to P^{\prime}\;\middle|\;\begin{aligned} &P,P^{\prime}\in\mathscr{V}^{\prime},\ \theta^{-1}(P^{\prime})\cap P\neq\emptyset\ \text{and there exist }R,R^{\prime}\in\mathscr{R},\ \text{such that}\\ &\overline{R}\cap(P\times M),\overline{R^{\prime}}\cap(P^{\prime}\cap M)\neq\emptyset\ \text{and }\Theta^{-1}(R^{\prime})\cap R\neq\emptyset\end{aligned}\right\}.$
@@
Since $\mathcal{X}$ is a Borel homoclinic class, recall that $\Sigma$ is an irreducible topological Markov shift. Therefore, there exists $\underline{y}\in\Sigma$ such that $\Sigma=\overline{\{\sigma^{i}\left(\underline{y}\right)\}_{i\in\mathbb{Z}}}.$ Take $\underline{x}\in\Sigma_{s}$ such that $\mathrm{proj}_{\Sigma}(\underline{x})=\underline{y}$. Let $\mathscr{G}^{\prime}_{s}\subset\mathscr{G}_{s}$ the maximal irreducible component of $\mathscr{G}_{s}$ containing $x_{0}$ (the 0-th element of $\underline{x}$) and define